OpenStudy (anonymous):

find coordinates of x/ y intercept and max or mini point on the grapg of these quad fuction. show answer by obtaining reasoing with sumbolic forms help: 9a ) f(x)=x(7-x) d) f(x)=-(x-3)(x+5) g) f(x)=2(x-3)(x-8)

OpenStudy (anonymous):

you at least have the x,y intercepts?

OpenStudy (anonymous):

i got nothing i got no idea how to do this

OpenStudy (anonymous):

let's find those first.

OpenStudy (anonymous):

ok

OpenStudy (apoorvk):

make x=0 and the value of y will give y-intercept. conversely true for x-intercept. for maxima or minima, differentiate and equate the differential to 0. the values of x for which it's zero, gives point of maxima and minima!

OpenStudy (anonymous):

write it out pelase

OpenStudy (anonymous):

finding the y-intercept is easy. make x=0 and evaluate each function.

OpenStudy (phi):

we went through the 2nd equation in quite a bit of detail.

OpenStudy (anonymous):

so for a it would be written f(0)=0(7-0)

OpenStudy (anonymous):

phi internet stopped working so lost everything we did

OpenStudy (anonymous):

ok, do the same for the other two...

OpenStudy (anonymous):

forget that. let's just work on one function first.

OpenStudy (anonymous):

oaky

OpenStudy (anonymous):

this quadratic function opens up so it will have a.... (max/min)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you answer? max or min?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

y = x(7-x) = 7x -7x^2. the coefficient of the x^2 term will tell you if it's opening upward or downward. If the number is positive, it opens upward. if negative, then downward.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

does it open upward or downard?

OpenStudy (anonymous):

up?

OpenStudy (phi):

This is going to be a frown shape. So it will have a maximum peak.

OpenStudy (anonymous):

ok

OpenStudy (phi):

Can you find the y intercept? f(0)=0(7-0)

OpenStudy (anonymous):

would it be 0

OpenStudy (phi):

That means multiply 0 times 7

OpenStudy (anonymous):

its o

OpenStudy (phi):

so you have the y-intercept. Now find the x-intercepts f(x)=x(7-x) what x values make this 0? x(7-x)=0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

and 7

OpenStudy (phi):

ok you have the y-intercept = 0, you have 2 x-intercepts 0 and 7 now find the x where the parabola peaks. It is half way between the 2 x intercepts add them together and divide by 2

OpenStudy (anonymous):

3.5

OpenStudy (anonymous):

dear take the second derivative of this function must be equal to zero, then if you get the positive answer then its minima and if you get the negative one then its maxima

OpenStudy (phi):

now to finish, find what the f(3.5) value is: f(x)=x(7-x) f(3.5) = 3.5*(7-3.5)

OpenStudy (anonymous):

how to solve that

OpenStudy (phi):

first do (7-3.5)=?

OpenStudy (anonymous):

3.5

OpenStudy (phi):

now multiply 3.5*(7-3.5) = 3.5 * 3.5 =

OpenStudy (anonymous):

igot a werid number

OpenStudy (anonymous):

is it 12.25

OpenStudy (anonymous):

?

OpenStudy (phi):